Which of the following statements is (are) correct? The shoeleather cost of inflation
(x) refers to the waste of resources used to maintain lower money holdings when inflation is high.
(y) should be approximately the same for an employed medical doctor and for an unemployed worker.
(z) is significant in countries with hyperinflation.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (x) only

You buy stock and its price rises at a rate of five percent. Inflation for the same period rises at a rate of five
percent. Before taxes you made
A. a nominal and real loss, but you pay taxes on the real loss.
B. a nominal and real gain, and you pay taxes on the nominal gain.
C. a nominal and real gain, but you pay taxes only on the real gain.
D. a nominal gain, but no real gain, so you pay no taxes on the capital gain.
E. a nominal gain, but no real gain, yet you pay taxes on the nominal gain.

Archie purchases some land for $30,000. He maintains it, but makes no improvements to it. One year later
he sells it for $32,000. Betty puts $30,000 in a savings account that pays 6% interest. Archie has to pay a
50% capital gains tax, Betty is in the 30% tax bracket. The inflation rate was 2%. Who had the higher
before-tax real gain and who had the higher after-tax real gain?
A. Archie had both the higher before-tax real gain and the higher after-tax real gain.
B. Archie had the higher before-tax real gain but Betty had the higher after-tax real gain.
C. Betty had the higher before-tax real gain but Archie had the higher after-tax real gain.
D. Betty had both the higher before-tax real gain and the higher after-tax real gain.
E. Betty had the higher before-tax real gain but they had the same after-tax real gain
Answer #1

The correct option to the first question is option c that is only X and z is true. The shoe leather cost of inflation is the lost of resources used to maintain the lower money holdings in the times of high inflation so by definition it is X is true. And also when there is hyperinflation the changes in the prices in the economy is very high and rapid. So people have to rush to banks to deposit their money holding and in times of hyperinflation such trips tens to be higher because prices changes very rapidly. And let's see why statement y is not true. An employed doctor will have much more resources as compared to an unemployed worker. And since shoe leather cost of inflation is based on loss of resources and therefore those who have more resource will have more show leather cost.

The correct option for second question is option E. Since the price rises by 5% and also the inflation rate is 5% so there is no real gain. And the nominal gain is 5% because the prices have risen by 5% so the nominal capital gain is 5%. And since there is nominal gain there will be some taxes on capital gains. So you see the correct option will be option E. That is a nominal gain, but no real gain, but you pay taxes on nominal gains.

The correct option for the third question is option B. That is Archie had higher before tax real gain but Betty had higher after tax real gain. The point is to look at the tax rate for the girls. Before tax gain for Archie is $2,000 and for Betty it is at 30,000/100 ×6 = $1,800. So Archie has higher before tax real gain. And after paying taxes the real capital gains for Archie will be 2000-1000 = $1,000 because the tax rate is 50% on capital gains. And after tax real gain for Betty is 1800/100×30 =$540 is the total tax on gains for Betty and the after tax gain is 1,800-540= $1,260. We can deduct 2% inflation to calculate real gain since rate of inflation is constant for both it doesn't matter for the purpose of comparing.